2. August 2004. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), EB = Erziehungswiss. Bibl., FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB = Informatikbibliothek, IB Nr. = Institutsbibliothek Nr., LB = HB– Lehrbuchsammlung, LS = HB–Lesesaal

Literatur über Boolesche Algebren und Boolesche Verbände

[1] H. Andraschko. Spezielle Boolesche Funktionen – Ein Weg zum Satz von Stone. Didaktik der Mathematik, 2(2):114–126, 1974. HB: Z 5339.

[2] Ladislav Beran. Über die Charakterisierung von Boole-Verbänden. Praxis der Mathematik, 17:98–103, 1975. MB: Z 101.

[3] C. Brink and J. Pretorius. Boolean circulants, groups, and relation algebras. The American Mathematical Monthly, 99(2):146–152, 1992.

[4] P. Dwinger. Introduction to Boolean Algebras. Physica–Verlag, Würzburg, 1971.

[5] George Grätzer. Lattice Theory: first concepts and distributive lattices. Freeman, 1971. MB: 6633.

[6] S. Koppelberg. Part I. General Theory of Boolean Algebras, pages 1–311. North–Holland, Amsterdam etc., 1989. MB: 14882 a.

[7] Mark V. Lawson. Finite Automata. Boca Raton: Chapman & Hall/CRC, 2004. ISBN 1-58488-255-7. HBZ. Algebraic automota theory, accessible at the undergraduate level, balanced to serve both mathematics and computer scienceoriented readers. Boolean operations, the syntactic monoid, algebraic language theory, graphs.

[8] Rudolf Lidl and Günter Pilz. Angewandte abstrakte Algebra. Vol. 1. Mannheim: BI, 1982. ISBN 3-411-01620-5. HB: BB1544-1+1.

[9] Rudolf Lidl and Günter Pilz. Angewandte abstrakte Algebra. Vol. 2. Mannheim: BI, 1982. ISBN 3-411- 3-411. HB: BB1544-2+1.

[10] Rudolf Lidl and Günter Pilz. Applied Abstract Algebra. Undergraduate texts in mathematics. New York [u.a.]: Springer, 1984, ^{2}1998. ISBN 0-387-96035-X, 3-540-96035-X; ^{2}0−387−98290−6. MB: 12609. Ch. 1: Lattices [§2 Boolean Algebras]. Ch. 2: Applications of Lattices [§1 Switching Circuits, §2 Propositional Logic]. Ch. 3: Finite Fields and Polynomials. Ch. 4: Coding Theory. Ch. 5.5: Pólya’s theory of enumeration. Ch. 6: Automata.

[11] Aleksandr Georgievic Pinus. Boolean Constructions in Universal Algebras. Mathematics and its Applications Vol. 242. Dordrecht: Kluwer, 1993. ISBN 0-7923-2117-0. HB: BB1469-242+1, MB: 16787.

[12] Egon Pracht. Algebra der Verbände. UTB 958. Paderborn: Schöningh, 1980. ISBN 3-506-99236-8. HB: Bf7312.

[13] Dieter Rüthing. Ein reduziertes Axiomensystem für Boolesche Verbände. Praxis der Mathematik, 17:158–165, 1975. MB: Z 101. Gemeint sind Boolesche Algebren: statt ∪, ∩, ’, 0, 1 werden nur ∪, ’ angesetzt.

[14] D. A. Vladimirov. Boolesche Algebren. Mathematische Lehrbücher und Monographien, Abt. 2, Mathematische Monographien 29. Berlin: Adad.-Verlag, 1972. MB: 6609.